OpenStudy (anonymous):

Solve by elimination x + 2y = 7 3x- 2y = -3

OpenStudy (anonymous):

@robz8

OpenStudy (anonymous):

same thing as before, except the y coefficients are already the same for you

OpenStudy (anonymous):

so just add the equations together x+2y=7 +3x-2y=-3 =4x = 4 x = 1 just like before, plug x back into one of the equations and solve for the value of y (1) + 2y = 7

OpenStudy (anonymous):

thx again

OpenStudy (nincompoop):